0.2w-0.47=0.53-0.2(2w-13)
Answer by checkley77(12844)

0.2w-0.47=0.53-0.2(2w-13)
.2w-.47=.53-.4w+2.6
.2w+.4w=.53+2.6+.47
.6w=3.6
w=3.6/.6
w=6 ans.
Proof:
.2*6-.47=.53-.2(2*6-13)
1.2-.47=.53-.2(12-13)
.73=.53-.2*-1
.73=.53+.2
.73=.73